# Some new findings concerning value distribution of a pair of delay-differential polynomials

### v1

@@ Abstract.
The paired Hayman’s conjecture of different types are considered. More accurately speaking, the zeros of a pair of $f^{n}(z)L(z,g)-a_{1}(z)$ and $g^{m}(z)L(z,f)-a_{2}(z)$ are characterized using different methods from those previously employed, where $f(z)$ and $g(z)$ are both transcendental entire functions, $L(z,f)$ and $L(z,g)$ are non-zero linear delay-differential polynomials, $\min\{n,m\}\geq 2$, $a_{1}(z),a_{2}(z)$ are non-zero small functions with relative to $f$ and $g$, or to $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$, respectively. These results give answers to three open questions raised by Gao, Liu[7] and Liu, Liu[15].
@@ 1. Introduction
Let $f$ be a meromorphic function in the complex plane $\mathbb{C}$. Assume that the reader is familiar with the standard notation and basic results of Nevanlinna theory, such as $m(r,f),~{}N(r,f)$,${}T(r,f)$, see [8] for more details. A meromorphic function $g$ is said to be a small function of $f$ if $T(r,g)=S(r,f)$, where $S(r,f)$ denotes any quantity that satisfies $S(r,f)=o(T(r,f))$ as $r$ tends to infinity, outside of a possible exceptional set of finite linear measure. We use $\rho(f)=\underset{r\rightarrow\infty}{\lim\sup}\frac{\log^{+}T(r,f)}{\log r}$ and $\rho_{2}(f)=\underset{r\rightarrow\infty}{\lim\sup}\frac{\log^{+}\log^{+}T(r,f)}{\log r}$ to denote the order and the hyper-order of $f$, respectively. In general, the linear delay-differential polynomial of $f$ is defined by
@@
Let $f(z)$ be a meromorphic function. If $f(z)-a(z)$ has finitely many zeros, then meromorphic $a(z)$ is called a generalized Picard exceptional function of $f(z)$. Furthermore, if $a(z)$ is also a small function of $f$, then we say $f$ has a generalized Picard exceptional small function $a(z)$. If $a(z)$ is a constant, then we say that $f$ has a generalized Picard exceptional value.
The well-known Picard theorem can be derived from the Nevanlinna’s second fundamental theorem. Here, we present the second fundamental theorem concerning three small functions[8, Theorem 2.5]: Let $f$ be a transcendental meromorphic, and let $a_{j}$ $(j=1,2,3)$ be a small function of $f$, then
@@
The generalizations of Nevanlinna’s second main theorem concerning $q\geq 3$ small functions was given by Yamanoi[21], where $q$ is an integer. It is also easy to see a transcendental meromorphic function has at most two generalized Picard exceptional small functions.
@@
[9, Theorem 10] If $f(z)$ is a transcendental entire function and $n\geq 2$ is a positive integer, then $f^{n}f^{\prime}$ can not have any non-zero generalized Picard exceptional values.
@@
[9] If $f(z)$ is a transcendental meromorphic function and $n$ is a positive integer, then $f^{n}f^{\prime}$ can not have any non-zero generalized Picard exceptional values.
@@
[10] If $f(z)$ is a transcendental entire function of finite order and $n\geq 2$ is a positive integer, then $f^{n}f(z+c)$ can not have any non-zero generalized Picard exceptional values, where $c$ is a non-zero constant.
For other results concerning the generalized Picard exceptional values of differences of meromorphic functions, please refer to [3, 6, 14, 13]. The generalized Picard exceptional small functions of delay-differential polynomials related to Hayman’s conjecture was considered by Liu, Liu and Zhou[16] using the ideas of common zeros and common poles, they obtained the following result.
@@
[16] Let $f(z)$ be a transcendental meromorphic function of hyper-order $\rho_{2}(f)<1$. If $n\geq 2k+6$ or if $n\geq 3$ and $f(z)$ is a transcendental entire function, then $f^{n}(z)f^{(k)}(z+c)$ can not have any non-zero generalized Picard exceptional small function $a(z)$ with respect to $f(z)$.
The condition $n\geq 2k+6$ in Theorem C has been weakened by Laine et al.[11] to $n>k+4$.
2. common generalized Picard exceptional small functions
In 2022, Gao and Liu considered the common generalized Picard exceptional small functions of a pair of complex delay-differential polynomials. These two delay-differential polynomials are the crossed variants of the complex differential polynomials in Hayman’s conjecture.
Theorem A .
[7] Let $f(z)$ and $g(z)$ be both transcendental meromorphic functions, $n$ and $k$ be non-positive integers, $a(z)$ be a non-zero small function with respect to $f$ and $g$. If one of the following conditions is satisfied:
@@
    $f$ and $g$ are both entire functions, $n\geq 3$, $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(h)=h^{(k)}(z+c)$ or L(h)=h(z+c)-h(z);
@@
    $f$ and $g$ are meromorphic functions, $\rho_{2}(f)<1$, and $\rho_{2}(g)<1$, $L(h)=h(z+c)$, $n\geq 4$ or $L(h)=h(z+c)-h(z)$, $n\geq 5$ or $L(h)=h^{(k)}(z+c)$, $n\geq k+4$, respectively;
@@
    $f$ and $g$ are entire functions, $n\geq 3$, $L(h)=h^{(k)}(z)$, or $f$ and $g$ are meromorphic functions, $n\geq k+4$, $L(h)=h^{(k)}(z)$, respectively;
then $f^{n}(z)L(g)$ and $g^{n}(z)L(f)$ cannot have a common non-zero generalized Picard exceptional small function $a(z)$.
Obviously, if $f\equiv g$, then Theorem A may reduce to Hayman’s conjecture of different types. Regarding Theorem A, Gao and Liu posed the following question for further investigation.
@@
[7, Question 1] Can we reduce $n\geq 3$ to $n\geq 2$ in Theorem A for entire functions $f$ and $g$? And what is the sharp value $n$ for meromorphic functions $f,g$?
Firstly, we give a positive answer to the first part of Question 1.
@@
Let $f(z)$ and $g(z)$ be both transcendental entire functions, $\min\{n,m\}\geq 2$, $a(z)$ be a non-zero small function with respect to $f$ and $g$. Then, $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $a(z)$, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small functions of $f$ and $g$ as its coefficients.
@@
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are linear delay-differential polynomial with small functions of $f$ and $g$ as their coefficients, $a(z)$ is a non-zero small funtion with respect to $f(z)$ and $g(z)$, then $f^{n}(z)L(g)$ and $g^{n}(z)L(f)$ can not have a common non-zero generalized Picard exceptional small function $a(z)$ by using the similar proof of Theorem 2.1. Combining this result with Theorem 2.1, we give a positive answer to Question 1 for entire functions $f$ and $g$.
@@
Obviously, if $f\equiv g$ is an entire function, Theorem 2.1 can be reduced to Theorems A and [12, Corollary 1.7], and improves Theorems B, C. What’s more, our approaches differ significantly from the methods outlined in Theorems A, B, C and A.
@@
If $m=n=1$, then Theorem 2.1 does not holds. For example, let $f(z)=e^{-p(z)}$, $g(z)=e^{p(z)}$, where $p(z)$ is a non-constant polynomial, then $fg-2=-1$ has no zeros.
From Theorem A, it is evident that a small function $a(z)$ respect to $f$ and $g$ can not serve as a common generalized Picard exceptional function for both $f^{n}(z)L(g)$ and $g^{n}(z)L(f)$. Naturally, this prompts the question[15, Question 1.1]: Given any two transcendental meromorphic functions $F(z)$ and $G(z)$, what can we obtain for their common or different generalized Picard exceptional values or small functions of $F$ and $G$. The direct consideration of this question is rather difficult, thus Liu and Liu[15] raised the following relatively simple question.
@@
Liu and Liu considered Question 2 and obtained partial results.
Theorem B .
@@
    If $\min\{m,n\}\geq 2$, then $f(z)^{n}g(z)$ and $g(z)^{m}f(z)$ can not have simultaneously non-zero Picard exceptional value.
@@
    If $n\geq 2$ and $m=1$ then $f(z)^{n}g(z)$ and $g(z)f(z)$ can not have simultaneously non-zero Picard exceptional value except that $f(z)=se^{T}(z)$ and $g(z)=s^{-2}a_{1}e^{-2T(z)}+a_{2}s^{-1}e^{-T(z)}$, where $T(z)$ is an entire function and $s$ is a complex constant. In this case, $a_{1}$ is the Picard exceptional value of $f(z)^{2}g(z)$ and $a_{2}$ is the Picard exceptional value of $f(z)g(z)$.
@@
    If $\min\{m,n\}\geq 3$, $k$ is a positive integer, then $f(z)^{n}g^{(k)}(z)$ and $g(z)^{m}f^{(k)}(z)$ can not have simultaneously non-zero Picard exceptional small functions.
Liu and Liu asked the following Question for Theorem B-$(3)$.
@@
[15, Remark 3.2] Can we reduce $n\geq 3$ to $n\geq 2$ in Theorem B-$(3)$?
We considered Questions 2 and 3 and obtained the following results.
@@
Let $f(z)$ and $g(z)$ be both transcendental entire functions.
@@
    If $\min\{m,n\}\geq 2$, then $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$ can not have simultaneously non-zero Picard exceptional small functions, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small functions of $f$ and $g$ as its coefficients.
@@
    If $n\geq 2$, $f^{n}(z)g(z)$ and $g(z)f(z)$ have simultaneously non-zero Picard exceptional small functions, then $f(z)=A(z)e^{B(z)}$ , $g(z)=A_{4}(z)e^{-nB(z)}+A_{3}(z)e^{-B(z)}$, where $A,A_{3},A_{4}$ are small entire function of $f$ and $g$, $B$ is an entire function. What’s more, $A^{n}A_{4}$ is the Picard exceptional small function of $f^{n}g$, $AA_{3}$ is the Picard exceptional small function of $fg$.
@@
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are linear delay-differential polynomial with small functions of $f$ and $g$ as its coefficients, then Theorem 2.2-$(1)$ still holds, using the similar proof of Theorem 2.2.
@@
We have found that $f(z)=e^{z}$ and $g(z)=2e^{-3z}+e^{-z}$ satisfy the relationships $f^{3}(z)g(z)=2+e^{-2z}$ and $g(z)f(z)=2e^{-2z}+1$. Consequently, $2$ is a Picard exceptional value of $f^{3}(z)g(z)$, $1$ is a Picard exceptional value of $g(z)f(z)$. However, these functions do not conform to the form $f(z)=se^{T}(z)$ and $g(z)=s^{-2}a_{1}e^{-2T(z)}+a_{2}s^{-1}e^{-T(z)}$ specified in Theorem B-$(2)$. This suggests that there may be a gap in the proof of Theorem B-$(2)$. Our Theorem 2.2-$(2)$ rectify the gap using different methods from those in Theorem B-$(2)$.
@@
We can provide an example for Theorem2.2-$(2)$ in the context of small functions: $f(z)=e^{z^{2}+z}e^{z^{5}}$, $g(z)=e^{-5z^{5}}+e^{z^{3}}e^{-z^{5}}$, which satisfies $f^{5}g=e^{5z^{2}+5z}+e^{-4z^{5}}e^{5z^{2}+5z+z^{3}}$, $fg=e^{z^{2}+z}e^{-4z^{5}}+e^{z^{2}+z+z^{3}}$. Here, $e^{5z^{2}+5z}$ is a Picard exceptional small function of $f^{5}g$, $e^{z^{2}+z+z^{3}}$ is a Picard exceptional small function of $fg$.
@@
Below, without loss of generality, we will prove Theorem 2.1 in two cases: whether $g$ is a small function of $f$ or not.
@@
$g$ is a small function of $f$. Since $a$ is small function of $f$, then $a$ is also a small function of $f^{n}L(z,g)$. We claim $f^{n}L(z,g)-a$ has infinite zeros. Otherwise, if $f^{n}L(z,g)-a$ has finite zeros, by the Weierstrass’s factorization theorem[18, pp. 145], we have
(3.1) $f^{n}(z)L(z,g)-a(z)=p(z)e^{b(z)},$
where $p$ is the infinite products formed by the zeros and poles of $f^{n}(z)L(z,g)-a(z)$, $b$ is an entire function, and they satisfy $T(r,p)=S(r,f^{n}L(z,g))=S(r,f)$ and $T(r,e^{b})=T(r,f^{n}L(z,g)+S(r,f)=T(r,f)+S(r,f)$.
@@
(3.2) $nf^{n-1}f^{\prime}L(z,g)+f^{n}L^{\prime}(z,g)-a^{\prime}=p_{1}e^{b},$
where $p_{1}=p^{\prime}+b^{\prime}p$. We claim $p^{\prime}+b^{\prime}p\not\equiv 0$. Otherwise $p^{\prime}+b^{\prime}p\equiv 0$, by By integrating, we can obtain $p(z)=c_{1}e^{-b(z)}$, where $c_{1}$ is a non-constant. Which is impossible, since $T(r,p)=S(r,f)$.
@@
(3.3) $p_{1}f^{n}L(z,g)-pnf^{n-1}f^{\prime}L(z,g)-pf^{n}L^{\prime}(z,g)=p_{1}a-pa^{\prime}.$
We claim $p_{1}a-pa^{\prime}\not\equiv 0$. Otherwise $p_{1}a-pa^{\prime}\equiv 0$, by By integrating, we can obtain $\frac{a}{p}=c_{2}e^{b}$, which is impossible, since $T(r,p)=S(r,f)$ and $T(r,a)=S(r,f)$.
Suppose $N(r,\frac{1}{f})\not=S(r,f)$, since coefficients of the equation (3.3) are small functions of $f$, then there exists a zero $z_{0}$ of $f$ such that the coefficients of the equation (3.3) are neither zero nor infinite at that point. Substituting $z_{0}$ into the equation (3.3), we easily obtain a contradiction. Therefore $N(r,\frac{1}{f})=S(r,f)$. By the The Weierstrass’s factorization theorem, we can let $f(z)=A(z)e^{B(z)}$, where $A$ and $B$ are entire function, such that $T(r,A)=S(r,f)$. By the second fundamental theorem concerning three small functions, we have
$T(r,f^{n})<\overline{N}(r,\frac{1}{f^{n}})+\overline{N}(r,\frac{1}{f^{n}-\frac{a}{L(z,g)}})+\overline{N}(r,f^{n})+S(r,f)=S(r,f),$
which is impossible. Therefore $f^{n}L(z,g)-a$ has infinite zeros. This means that $a(z)$ cannot be a common non-zero generalized Picard exceptional for both $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$.
@@
We only need to consider $m(r,g)\leq m(r,f)$, since the case $m(r,f)\leq m(r,g)$ can be studied by the same methods. Below, we will consider two cases: $m(r,g)\leq m(r,f)$ and whether $a$ is a small function of $f_{2}$ or not.
@@
$m(r,g)\leq m(r,f)$ and $a$ is a small function of $f_{2}$. Since $m(r,g)\leq m(r,f)$, then from (2), we get $m(r,f_{1})\geq(n-1)m(r,f)$, which means $a(z)$ is a small function of $f_{1}$. We claim at least one of $f^{n}L(z,g)-a$ and $g^{m}L(z,f)-a$ has infinite zeros. Otherwise, suppose $f^{n}L(z,g)-a$ and $g^{m}L(z,f)-a$ has finite zeros. Since $a$ is a small function of $f_{1}$ and $f_{2}$, by the Weierstrass’s factorization theorem, we have
(3.6) $f^{n}L(z,g)-a=p_{1}e^{b_{1}(z)},$

### v3

@@ Abstract.
The paired Hayman’s conjecture of different types are considered. More accurately speaking, the zeros of a pair of $f^{n}L(z,g)-a_{1}(z)$ and $g^{m}L(z,f)-a_{2}(z)$ are characterized using different methods from those previously employed, where $f$ and $g$ are both transcendental entire functions, $L(z,f)$ and $L(z,g)$ are non-zero linear delay-differential polynomials, $\min\{n,m\}\geq 2$, $a_{1},a_{2}$ are non-zero small functions with relative to $f$ and $g$, or to $f^{n}(z)L(z,g)$ and $g^{m}(z)L(z,f)$, respectively. These results give answers to three open questions raised by Gao, Liu[7] and Liu, Liu[15].
@@ 1. Introduction
Let $f$ be a meromorphic function in the complex plane $\mathbb{C}$. Assume that the reader is familiar with the standard notation and basic results of Nevanlinna theory, such as $m(r,f),~{}N(r,f)$,${}T(r,f)$, see [8] for more details. A meromorphic function $g$ is said to be a small function of $f$ if $T(r,g)=S(r,f)$, where $S(r,f)$ denotes any quantity that satisfies $S(r,f)=o(T(r,f))$ as $r$ tends to infinity, outside of a possible exceptional set of finite linear measure. $\rho(f)=\underset{r\rightarrow\infty}{\lim\sup}\frac{\log^{+}T(r,f)}{\log r}$ and $\rho_{2}(f)=\underset{r\rightarrow\infty}{\lim\sup}\frac{\log^{+}\log^{+}T(r,f)}{\log r}$ are used to denote the order and the hyper-order of $f$, respectively. In general, the linear delay-differential polynomial of $f$ is defined by
@@
Let $f$ be a meromorphic function. If $f-\alpha$ has finitely many zeros, then meromorphic function $\alpha$ is called a generalized Picard exceptional function of $f$. Furthermore, if $\alpha$ is also a small function of $f$, then we say $f$ has a generalized Picard exceptional small function $\alpha$. If $\alpha$ is a finite constant, then we say that $f$ has a generalized Picard exceptional value $\alpha$.
The well-known Picard theorem can be derived from the Nevanlinna’s second fundamental theorem. Here, we present the second fundamental theorem concerning three small functions[8, Theorem 2.5]: Let $f$ be a transcendental meromorphic function, and let $a_{j}$ $(j=1,2,3)$ be small functions of $f$, then
@@
The generalizations of Nevanlinna’s second main theorem concerning $q~{}(\geq 3)$ small functions was given by Yamanoi[21], where $q$ is an integer. It is also easy to see a transcendental meromorphic function has at most two generalized Picard exceptional small functions.
@@
[9, Theorem 10] If $f$ is a transcendental entire function and $n~{}(\geq 2)$ is a positive integer, then $f^{n}f^{\prime}$ cannot have any non-zero generalized Picard exceptional values.
@@
[9] If $f$ is a transcendental meromorphic function and $n$ is a positive integer, then $f^{n}f^{\prime}$ cannot have any non-zero generalized Picard exceptional values.
@@
[10] If $f$ is a transcendental entire function of finite order and $n~{}(\geq 2)$ is a positive integer, then $f^{n}f(z+c)$ cannot have any non-zero generalized Picard exceptional values, where $c$ is a non-zero constant.
For other results concerning the generalized Picard exceptional values of differences of meromorphic functions, please refer to [3, 6, 13, 14]. The generalized Picard exceptional small functions of delay-differential polynomials related to Hayman’s conjecture was considered by Liu, Liu and Zhou[16], and they obtained the following result.
@@
[16] Let $f$ be a transcendental meromorphic function of hyper-order $\rho_{2}(f)<1$, and let $c$ be a finite constant. If $n\geq 2k+6$ or if $n\geq 3$ and $f$ is a transcendental entire function, then $f^{n}f^{(k)}(z+c)$ can not have any non-zero generalized Picard exceptional small function $a$ with respect to $f$.
Later, the condition $n\geq 2k+6$ in Theorem C has been weakened by Laine et al.[11] to $n>k+4$.
2. Hayman’s conjecture on a pair of delay-differential polynomials
In 2022, Gao and Liu[7] considered the Hayman’s conjecture regarding a pair of delay-differential polynomials.
Theorem D .
[7] Let $f$ and $g$ be both transcendental meromorphic functions, $n$ and $k$ be non-negative integers, $\alpha$ be a non-zero small function with respect to $f$ and $g$. If one of the following conditions is satisfied:
@@ 2. Hayman’s conjecture on a pair of delay-differential polynomials
    $f$ and $g$ are both entire functions, $n\geq 3$, $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,h)=h^{(k)}(z+c)$ or $L(z,h)=h(z+c)-h(z)$;
@@
    $f$ and $g$ are meromorphic functions, $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,h)=h(z+c)$, $n\geq 4$ or $L(z,h)=h(z+c)-h(z)$, $n\geq 5$ or $L(z,h)=h^{(k)}(z+c)$, $n\geq k+4$, respectively;
@@
    $f$ and $g$ are entire functions, $n\geq 3$, $L(z,h)=h^{(k)}(z)$, or $f$ and $g$ are meromorphic functions, $n\geq k+4$, $L(z,h)=h^{(k)}(z)$, respectively;
then $f^{n}L(z,g)$ and $g^{n}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$.
Obviously, if $f\equiv g$, then Theorem D may reduce to Hayman’s conjecture of different types. Regarding Theorem D, Gao and Liu posed the following question for further investigation.
@@
[7, Question 1] Can we reduce $n\geq 3$ to $n\geq 2$ in Theorem D for entire functions $f$ and $g$? And what is the sharp value $n$ for meromorphic functions $f,g$?
By considering a broader class of a pair of delay-differential polynomials, we give a positive answer to the first part of Question 1.
@@
Let $f$ and $g$ be both transcendental entire functions, $\min\{n,m\}\geq 2$, $\alpha$ be a non-zero small function with respect to $f$ and $g$. Then, $f^{n}L(z,g)$ and $g^{m}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small functions of $f$ and $g$ as its coefficients.
@@
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are non-zero linear delay-differential polynomial with small functions of $f$ and $g$ as their coefficients, $\alpha$ is a non-zero small funtion with respect to $f$ and $g$, then $f^{n}L(z,g)$ and $g^{n}L(z,f)$ cannot have a common non-zero generalized Picard exceptional small function $\alpha$ by using the similar proof of Theorem 2.1. Combining this result with Theorem 2.1, we give a positive answer to Question 1 for entire functions $f$ and $g$.
@@
Obviously, if $f\equiv g$ is an entire function, Theorem 2.1 can be reduced to Theorems A, B and [12, Corollary 1.7], and improves Theorems C. What’s more, our approaches differ significantly from the methods outlined in Theorems A, B, C and D.
@@
If $m=n=1$, then Theorem 2.1 does not holds. For example, let $f(z)=e^{-p(z)}$, $g(z)=e^{p(z)}$, where $p$ is a non-constant polynomial, then $fg-2=-1$ has no zeros.
From Theorem D, it is evident that a small function $\alpha$ respect to $f$ and $g$ can not serve as a common generalized Picard exceptional function for both $f^{n}L(,g)$ and $g^{n}L(z,f)$. Naturally, this prompts the question[15, Question 1.1]: Given any two transcendental meromorphic functions $F(z)$ and $G(z)$, what can we obtain for their common or different generalized Picard exceptional values or small functions of $F$ and $G$? The direct consideration of this question is rather difficult, thus Liu and Liu[15] raised the following question.
@@
Liu and Liu[15] considered Question 2 and obtained partial results.
Theorem E .
@@
    If $\min\{m,n\}\geq 2$, then $f^{n}g$ and $g^{m}f$ can not have simultaneously non-zero generalized Picard exceptional value.
@@
    If $n\geq 2$ then $f^{n}g$ and $gf$ can not have simultaneously non-zero generalized Picard exceptional value except that $f(z)=se^{T}(z)$ and $g(z)=s^{-2}a_{1}e^{-2T(z)}+a_{2}s^{-1}e^{-T(z)}$, where $T(z)$ is an entire function and $s$ is a complex constant. In this case, $a_{1}$ is the Picard exceptional value of $f^{2}g$ and $a_{2}$ is the Picard exceptional value of $fg$.
@@
    If $\min\{m,n\}\geq 3$, $k$ is a positive integer, then $f^{n}g^{(k)}$ and $g^{m}f^{(k)}$ can not have simultaneously non-zero Picard generalized exceptional small functions.
Liu and Liu asked the following question related Theorem E-$(3)$.
@@
[15, Remark 3.2] Can we reduce $n\geq 3$ to $n\geq 2$ in Theorem E-$(3)$?
We consider Questions 2 and 3 and obtain the following results.
@@
Let $f$ and $g$ be both transcendental entire functions.
@@
    If $\min\{m,n\}\geq 2$, then $f^{n}L(z,g)$ and $g^{m}(z)L(z,f)$ cannot have simultaneously non-zero generalized Picard exceptional small functions, where $L(z,f)$ and $L(z,g)$ are non-zero linear differential polynomial with small functions of $f$ and $g$ as its coefficients.
@@
    If $n\geq 2$, then $f^{n}g$ and $gf$ can not have simultaneously non-zero generalized Picard exceptional value except that $f(z)=A(z)e^{B(z)}$, $g(z)=A_{4}(z)e^{-nB(z)}+A_{3}(z)e^{-B(z)}$, where $A,A_{3},A_{4}$ are small entire function of $f$ and $g$, $B$ is an entire function. What’s more, $A^{n}A_{4}$ is the Picard exceptional small function of $f^{n}g$, $AA_{3}$ is the Picard exceptional small function of $fg$.
@@
If $\rho_{2}(f)<1$ and $\rho_{2}(g)<1$, $L(z,f)$ and $L(z,g)$ are linear delay-differential polynomial with small functions of $f$ and $g$ as its coefficients, then Theorem 2.2-$(1)$ still holds by using the similar proof of Theorem 2.2.
@@
Note that $f(z)=e^{z}$ and $g(z)=2e^{-3z}+e^{-z}$ satisfy the relationships $f^{3}(z)g(z)=2+e^{2z}$ and $g(z)f(z)=2e^{-2z}+1$. Consequently, $2$ is a Picard exceptional value of $f^{3}g$, $1$ is a Picard exceptional value of $gf$. However, $g(z)=2e^{-3z}+e^{-z}$ does not conform to the form $g(z)=s^{-2}a_{1}e^{-2T(z)}+a_{2}s^{-1}e^{-T(z)}$ specified in Theorem E-$(2)$. This suggests that there may be a gap in the proof of Theorem E-$(2)$. Our Theorem 2.2-$(2)$ rectify the gap using different methods from those in Theorem E.
@@
An example of Theorem 2.2-$(2)$ in the context of small functions is given as follows: $f(z)=e^{z^{2}+z}e^{z^{5}}$, $g(z)=e^{-5z^{5}}+e^{z^{3}}e^{-z^{5}}$, which satisfies $f^{5}g=e^{5z^{2}+5z}+e^{4z^{5}}e^{5z^{2}+5z+z^{3}}$, $fg=e^{z^{2}+z}e^{-4z^{5}}+e^{z^{2}+z+z^{3}}$. Here, $e^{5z^{2}+5z}$ is a Picard exceptional small function of $f^{5}g$, $e^{z^{2}+z+z^{3}}$ is a Picard exceptional small function of $fg$.
@@
Without loss of generality, we will prove Theorem 2.1 in two cases: whether $g$ is a small function of $f$ or not.
@@
$g$ is a small function of $f$. Since $\alpha$ is a small function of $f$, then $\alpha$ is also a small function of $f^{n}L(z,g)$. We claim $f^{n}L(z,g)-\alpha$ has infinitely many zeros. Otherwise, if $f^{n}L(z,g)-\alpha$ has finitely many zeros, by the Weierstrass’s factorization theorem[18, pp. 145], we have
(3.1) $f^{n}(z)L(z,g)-\alpha(z)=p(z)e^{b(z)},$
where $p$ is the infinite products formed by the zeros and poles of $f^{n}L(z,g)-\alpha$, $b$ is an entire function, and they satisfy $T(r,p)=S(r,f^{n}L(z,g))=S(r,f)$ and $T(r,e^{b})=T(r,f^{n}L(z,g)+S(r,f)=O(T(r,f))+S(r,f)$.
@@
(3.2) $nf^{n-1}f^{\prime}L(z,g)+f^{n}L^{\prime}(z,g)-\alpha^{\prime}=p_{1}e^{b},$
where $p_{1}=p^{\prime}+b^{\prime}p$. We claim $p^{\prime}+b^{\prime}p\not\equiv 0$. Otherwise $p^{\prime}+b^{\prime}p\equiv 0$, by integrating, we can obtain $p(z)=c_{1}e^{-b(z)}$, where $c_{1}$ is a non-constant. Which is impossible, since $T(r,p)=S(r,f)$.
@@
(3.3) $p_{1}f^{n}L(z,g)-pnf^{n-1}f^{\prime}L(z,g)-pf^{n}L^{\prime}(z,g)=p_{1}\alpha-p\alpha^{\prime}.$
We claim $p_{1}\alpha-p\alpha^{\prime}\not\equiv 0$. Otherwise $p_{1}\alpha-p\alpha^{\prime}\equiv 0$, by integrating, we can obtain $\frac{\alpha}{p}=c_{2}e^{b}$, which is impossible, since $T(r,p)=S(r,f)$ and $T(r,\alpha)=S(r,f)$.
Suppose $N(r,\frac{1}{f})\not=S(r,f)$, since coefficients of the equation (3.3) are small functions of $f$, then there exists a zero $z_{0}$ of $f$ such that the coefficients of the equation (3.3) are neither zero nor infinite at that point. Substituting $z_{0}$ into the equation (3.3), we easily obtain a contradiction. Therefore $N(r,\frac{1}{f})=S(r,f)$. By the Weierstrass’s factorization theorem, let $f(z)=A(z)e^{B(z)}$, where $A$ and $B$ are entire function, such that $T(r,A)=S(r,f)$. By the second fundamental theorem concerning three small functions and (3.1), we have
$T(r,f^{n})<\overline{N}(r,\frac{1}{f^{n}})+\overline{N}(r,\frac{1}{f^{n}-\frac{\alpha}{L(z,g)}})+\overline{N}(r,f^{n})+S(r,f)=S(r,f),$
which is impossible. Therefore $f^{n}L(z,g)-\alpha$ has infinitely many zeros. This means that $\alpha$ cannot be a common non-zero generalized Picard exceptional small function for both $f^{n}L(z,g)$ and $g^{m}L(z,f)$.
@@
We only need to consider $m(r,g)\leq m(r,f)$, since the case $m(r,f)\leq m(r,g)$ can be studied by the same methods. Below, we will consider two subcases: $m(r,g)\leq m(r,f)$ and whether $\alpha$ is a small function of $f_{2}$ or not.
@@
$m(r,g)\leq m(r,f)$ and $\alpha$ is a small function of $f_{2}$. Since $m(r,g)\leq m(r,f)$, then from (2), we get $(n+1)m(r,f)+S(r,f)\geq m(r,f_{1})\geq(n-1)m(r,f)$, which means $\alpha$ is a small function of $f_{1}$. We claim at least one of $f^{n}L(z,g)-\alpha$ and $g^{m}L(z,f)-\alpha$ has infinitely many zeros. Otherwise, suppose $f^{n}L(z,g)-\alpha$ and $g^{m}L(z,f)-\alpha$ have finitely many zeros. Since $\alpha$ is a small function of $f_{1}$ and $f_{2}$, by the Weierstrass’s factorization theorem, we have
(3.6) $f^{n}L(z,g)-\alpha=p_{1}e^{b_{1}(z)},$